### v2

# Getting Python Types Right with RightTyper

Abstract.
[…]
This paper presents RightTyper, a novel hybrid approach for Python that produces accurate and precise type annotations grounded in actual program behavior. RightTyper grounds inference in types observed during actual program execution and combines these observations with static analysis and name resolution to produce substantially higher-quality type annotations than prior approaches. Through principled, statistically guided adaptive sampling, RightTyper balances runtime overhead with the need to observe sufficient execution behavior to infer high-quality type annotations. We evaluate RightTyper against static, dynamic, and AI-based systems on both synthetic benchmarks and real-world code, and find that it consistently achieves higher semantic similarity to ground-truth and developer-written annotations, respectively, while incurring only approximately 25% runtime overhead.
Type Inference, Type Annotations, Python, Sampling, Dynamic Analysis, Static Analysis
footnotetext: Work done at Amazon Web Services.
1. Introduction
Python is now firmly established as one of the most popular programming languages. Despite being dynamically typed, Python has supported static type annotations since version 3.5, released in 2015 (pep484). While the Python runtime does not enforce these type annotations, separate tools such as mypy and pyright (mypy; pyright) use them to perform static type checking, catching bugs before deployment that would otherwise manifest as runtime errors. Beyond static verification, type annotations improve code readability, enable better IDE support (code completion and refactoring), and support runtime validation frameworks (e.g., pydantic and beartype). Collectively, these benefits can significantly reduce debugging time and improve software reliability.
Despite these many uses and a decade of language support, Python codebases remain largely unannotated. A recent study of almost ten thousand popular Python project repositories found that only 7% of them use type annotations at all. Among those with annotations, only about 8% of function arguments and return types are actually annotated (10.1145/3540250.3549114). This gap highlights the practical challenges developers face in manually adding annotations, especially to large codebases. To address this, type inference tools aim to automatically infer and insert type annotations, reducing manual effort and making static typing more accessible.
Past approaches to Python type inference fall into three broad categories: static, AI-based, and dynamic. Static approaches reason about code without executing it, using information such as program structure, operations, and control flow to infer types (static-program-analysis). However, they are limited by dynamic language features, either supporting only subsets of the language or leaving portions of the code unannotated (10.1145/1806596.1806598; 10.1007/978-3-642-03237-0_17). They also tend to be imprecise, as they conservatively over-approximate program behavior (ruby-dynamic). For example, if a function could theoretically operate on either an integer or a string, a static tool may reflect both in its inference, even if only one is ever actually used—diminishing the effectiveness of static checkers for catching bugs.
AI-based approaches operate on source code to estimate the likely types of program elements, most commonly by applying machine learning models. These methods are better equipped than static approaches to handle the complexities of dynamic languages, but they sacrifice soundness: predicted types may not reflect actual program behavior, undermining the effectiveness of type checking through false positives or false negatives. Some hybrid systems attempt to address this unsoundness by using static analysis to validate model predictions (typewriter; hityper). However, this validation can only discard incorrect type predictions, potentially leaving more of the code unannotated. Moreover, these methods struggle with rare types, and many support only a limited type vocabulary, making them ill-equipped to handle user-defined or evolving types (type4py).
[…]
1.1. Contributions
This paper proposes RightTyper, a novel hybrid type inference approach for Python. RightTyper sidesteps pitfalls of static and AI-based methods by sampling types observed during actual program execution. It employs adaptive sampling strategies to minimize runtime overhead while maintaining inference accuracy. RightTyper monitors function calls using a Poisson process, in which observations occur only during brief, randomly timed capture windows, allowing most instrumentation to be disabled outside these windows. For container types, RightTyper fully scans small containers, but applies Good–Turing estimation as a stopping criterion for larger ones, an estimator developed by Turing to estimate unseen events (good-turing; vanderWalt2021). Additionally, RightTyper tracks previously seen containers to avoid re-typing them, while continuing to sample to detect changes. Together, these strategies achieve low overhead (overall $\approx$25%) while using statistically principled sampling to obtain representative coverage.
By combining observed types with information from static analysis, type aggregation, and name resolution, RightTyper emits annotations that substantially improve upon those produced by prior approaches. This paper makes the following contributions:
  • It introduces RightTyper, a novel dynamic–static hybrid type inference approach for Python (§2);
  • It presents a prototype implementation of RightTyper, released as open-source software on GitHub (§3);
  • It presents an enhanced implementation of TypeSim—a semantic type similarity metric proposed by prior work (typybench)—that addresses various issues in the original implementation and is also released on GitHub (§4.1);
  • It empirically compares the types inferred by RightTyper to those from prior work, demonstrating a marked improvement in the quality of annotations (§4.2, §4.3);
  • It evaluates RightTyper’s runtime overhead against other inference tools, showing that it is significantly more efficient, and fast enough for practical use (§4.4).
2. Approach
RightTyper executes the target program using a combination of custom instrumentation and Python’s standard monitoring mechanisms. During execution, it inspects runtime objects—such as function parameters, return values, and variables—to observe their types, and samples the contents of container objects (e.g., lists and dictionaries) to infer their complete types. To manage overhead, RightTyper controls instrumentation using a Poisson process, enabling and disabling monitoring to create brief, randomly timed capture windows. After execution, RightTyper combines these runtime observations with types extracted via static analysis, identifies recurring type patterns, and simplifies them to produce high-quality annotations. Figure 1 provides an overview.
Figure 1. RightTyper Overview: RightTyper executes the target program under instrumentation, dynamically managing it to create Poisson-distributed observation windows (§2.1). As the instrumentation delivers execution events, it observes and collects types used at runtime (§2.2, §2.3). After execution, RightTyper combines these runtime observations with types extracted via static analysis to generate type annotations (§2.4–§2.8).
2.1. Instrumentation and Event Sampling
Most of RightTyper’s instrumentation utilizes sys.monitoring, a low-overhead, fine-grained mechanism introduced in Python 3.12 for tracking various events during execution (pep669). This feature allows selective activation and deactivation of monitoring for specific portions of user code. To capture events not supported by sys.monitoring, RightTyper additionally instruments the code (§3).
To minimize runtime overhead while obtaining representative type observations, RightTyper governs execution event monitoring using a Poisson process. Each function—or, more precisely, each code object, most of which correspond to functions—undergoes an initial warmup phase during which the first $k$ invocations (with default k=5k{=}5) are captured unconditionally. This warmup helps ensure sufficient observations for basic type inference, even for infrequently executed functions and very short executions.
After warmup, monitoring transitions to Poisson-timed sampling: a timer fires at exponentially distributed intervals (with configurable rate $\lambda$, default $\lambda{=}2\ \mathrm{Hz}$), enabling observation only during brief capture windows. This approach yields statistically principled coverage while maintaining low average overhead, as the majority of function calls execute without instrumentation. Together, the warmup phase and Poisson sampling ensure that RightTyper observes all executed functions and captures a representative sample of their invocations over time. RightTyper combines the types of a function’s arguments, variables, and return values—observed separately across multiple execution events—into a call trace, which it later uses to generate annotations. By capturing all types involved in a call, call traces enable RightTyper to identify recurring type patterns and produce more precise annotations (see §2.4).
2.2. Typing Objects
As it processes monitoring events, RightTyper examines the objects passed into and returned from each function, as well as any variables, to infer their types.
Figure 2. Capturing variable initializations: Although RightTyper primarily captures variable types at function exit, its loader also records constant-valued initializations. This technique enables correct typing of optionally typed variables, such as prev_type in this excerpt from the black source code.
Variables
RightTyper captures variable types, including object and class attributes, by inspecting their values when a function returns or unwinds. This design reflects a performance trade-off: in Python, variables may change type on any assignment, making continuous tracking expensive. Capturing values at function exit instead provides a practical balance between accuracy and overhead. Static analysis performed by RightTyper’s loader supports this process by identifying the variables and attributes that should be tracked so they can be queried efficiently at function exit. The loader also records any constants assigned to a variable when it is first introduced, which is particularly useful for correctly typing variables with optional values, as illustrated in Figure 2.
Type Names and Import Paths
Because most types—aside from a few built-ins—must be imported for use in annotations, RightTyper must determine the module and name under which each type is available; together, we refer to these here as the import path. Because Python’s introspection does not always provide usable import paths—especially for types implemented in C—RightTyper constructs a mapping from type objects to import paths, applying heuristics to identify a canonical import for each type. This approach yields more natural and consistent annotations across the codebase (§3). For commonly used built-in objects that lack importable type definitions, such as iterators, RightTyper falls back on structural subtyping, using protocols such as Iterator (pep484).
2.3. Typing Generics
When a type is generic, RightTyper also attempts to infer its type parameters. While some generic instances expose their parameters directly, many others—such as standard container types—do not, and often contain heterogeneous elements; exhaustively enumerating their contents becomes very expensive for large instances. RightTyper addresses this challenge through adaptive sampling, combining Good–Turing estimation with novelty detection to efficiently infer container element types.
Containers
RightTyper tracks containers and, upon first encountering a container or detecting a change, inspects its contents, distinguishing between small and large containers. RightTyper fully scans small containers (of size $\leq 32$ by default), while it analyzes large containers using probabilistic sampling with a Good–Turing stopping criterion. Originally developed during World War II by Alan Turing and later formalized by I. J. Good, the Good–Turing estimator estimates the probability of unseen events (good-turing; vanderWalt2021).
For large containers, after collecting a minimum number of samples (default 32), RightTyper computes the singleton ratio $\hat{r}$—the proportion of observed types that appear exactly once:
$\hat{r}=\frac{\sum_{t\in C}\mathbf{1}[C(t)=1]}{n}$
where $C$ is a counter mapping each observed type $t$ to its count $C(t)$, $n$ is the number of samples, and $\mathbf{1}[\cdot]$ is the indicator function. When this ratio falls below a threshold $\tau$ (default $\tau=0.05$) for all parameters (e.g., dictionary key and value types are counted separately), sampling terminates. A hard cap (default 64 samples) further prevents pathological cases from dominating execution time. RightTyper aggregates all types observed across inspections into a cumulative history, from which it generates annotations.
To handle containers that evolve over time, RightTyper tracks container size and employs probabilistic spot-checking. When a previously-sampled container is revisited, a re-scan (or re-sampling) occurs if the container’s size has changed, or a random spot-check discovers a novel type not present in the cumulative history.
Tracking containers is not only important for performance, but also for correctness: mutable containers are typically invariant with respect to their type parameters. For example, if RightTyper observes a function taking a list[int] and later a list[str], these observations should lead to a list[int|str] type only if they correspond to the same container instance; otherwise, they reflect containers with different element types and RightTyper should not merge them, since a list[int] does not permit adding str elements.
Iterators
Iterators pose special challenges for typing because their interface does not allow inspection of return values without destructively updating their state. For iterators implemented in Python, whose iteration method calls are visible through sys.monitoring, RightTyper begins with a partial type and updates it to a complete one once it observes an iteration. Built-in iterators, however, do not trigger monitoring events and require special handling (§3).
Numerical Arrays
Numerical array shape mismatches are a common source of errors in machine learning programs (gradualtensorshapechecking). While numpy does not support array shape annotations at the time of writing, RightTyper can optionally include shape information using the format that jaxtyping supports, a package that enables runtime shape checking. For instance, RightTyper would annotate a $2\times 3$ numpy array as Float64[ndarray, "2 3"]. When emitting such annotations for arrays, RightTyper also attempts to identify patterns in the shapes observed across traces, replacing them with variables. Figure 3 shows an example.
Figure 3. Annotation from shape pattern: As it does for type patterns, when generating jaxtyping-style annotations, RightTyper identifies recurring patterns in the observed array shapes across traces (top section) and replaces them with variables (bottom section), in a manner inspired by Hindley–Milner–style type generalization. To the best of our knowledge, no other existing typing tools support inferring and annotating array dimensions in this way.
2.4. Typing Functions
Figure 4. Naïve typing with unions: This example illustrates a function with interdependent argument and return types. The union of observed types (bottom), emitted by MonkeyType, is overly permissive and incorrectly allows mixed-type inputs that result in a type error.
After the target program finishes executing, RightTyper uses the collected call traces to generate type annotations for each function. When the type of an argument or return value varies across traces, the simplest annotation is a union of the observed types. For instance, if a function receives a number argument in one sample and a string in another, it may be annotated as f(a: float|str). While unions offer a flexible way to describe multiple observed types, they can also be overly permissive, potentially leading to errors or false negatives during static analysis. In particular, they may fail to capture essential relationships between arguments. For example, Figure 4 illustrates the function add, which returns the sum of two numbers or the concatenation of two strings. Naïvely annotating both arguments as float|str permits invalid combinations – such as one argument being a float and the other a string – that would result in either a static type error or, if undetected, a runtime error. In this example, because the arguments are used directly with the addition operator, mypy can detect the error, issuing the message shown in Figure 4.
For this reason, rather than immediately constructing unions, RightTyper first searches for patterns in the observed types. When it detects consistent variability across argument or return value types, RightTyper introduces a type argument to capture this shared variability, binding it to the concrete types encountered during execution. The center portion of Figure 5 illustrates the resulting annotation. The search operates recursively, allowing it to detect patterns nested within type arguments. For instance, a function that returns one of the elements of a list might be annotated as f(l: list[T]) -> T. Algorithm 1 describes the process in more detail. If the target Python version does not support type arguments, RightTyper instead defines a type variable to achieve the same effect, albeit in a more verbose manner, as the bottom section of Figure 5 shows.
Algorithm 1 Call trace generalization: The algorithm examines the types observed at each position (argument or return value) in a collection of call traces, generating a type for each corresponding position in the function signature. It transposes the call traces to group types by position and operates recursively, enabling it to identify patterns in the arguments of generic types. The symbol $\mathbin{\oplus}$ denotes concatenation.
function Generalize(tracestraces)
   function Rebuild(pos_typespos\_types)
    if all pos_typespos\_types are specializations of a generic $G$ then
      $arg\_positions\leftarrowtranspose$(arguments(pos_typespos\_types))
      $arguments\leftarrow[\ ]$
      for each $p$ in arg_positionsarg\_positions do
       $arguments\leftarrow arguments\ \mathbin{\oplus}$ Rebuild($p$)
      return $G[arguments]$
    else if pos_typespos\_types occurs more than once then
      $T\leftarrow$ assign or retrieve typevar for pos_typespos\_types
      return $T$
    else
      return Union(pos_typespos\_types)
   $signature\leftarrow[\ ]$
   for each $p$ in transpose(tracestraces) do
    $signature\leftarrow signature\ \mathbin{\oplus}$ Rebuild($p$)
   return signaturesignature
Figure 5. RightTyper recognizes type patterns: recognizing that add (Figure 4) is consistently called with either strings or numbers, RightTyper annotates the function using either a type argument (if supported by the target Python version; top portion) or a generated type variable (rt_T1; bottom portion).
2.5. Typing Methods
Methods require special handling because their arguments and return values may reference the current instance, which could be a subtype of the defining class. Consider Figure 6: when scale() is invoked on a Line object, annotating self: Line would be incorrect, as the mypy error in the figure demonstrates. Instead, RightTyper annotates using the defining class, or Self from the typing module when targeting Python versions that support it (pep673).
Figure 6. Inheritance complicates typing: in the top section, class Line inherits the scale() method from its parent class, Shape. When scale() is invoked on a Line object, its self parameter refers to that object—an instance of Line. However, annotating self: Line (bottom section), as MonkeyType and PyAnnotate do, is incorrect, as demonstrated by the mypy error.
Figure 7. Method overriding complicates typing: In the top section, class Line overrides the draw() method inherited from its parent class, Shape. Although Line.draw() is only ever invoked with Screen as its medium, the overriding method must remain compatible with all argument types accepted by Shape.draw(). Accordingly, RightTyper annotates the parameter using the union of the observed type(s) and those accepted by the parent method (bottom section). MonkeyType and PyAnnotate only annotate using the types they observe, which leads to errors (not shown).
Method arguments also require handling to avoid violating the Liskov Substitution Principle (LSP) (10.1145/62138.62141). Consider Figure 7, where Line overrides draw from Shape. Even though callers only invoke Line.draw with Screen as the medium argument, annotating it as medium: Screen would violate the LSP by excluding other valid media accepted by the base class. To address this, RightTyper inspects parent classes to determine whether a method overrides one from a superclass, and if so, annotates the overriding method’s arguments using the union of observed and inherited types.
To determine parent class annotations, RightTyper consults type stubs from typeshed when necessary, since the vast majority of the standard library lacks inline type annotations (typeshed).
2.6. Simplifying Types
If RightTyper cannot identify a consistent pattern across function argument or return value types (§2.4), or when typing variables, it constructs a union of the observed types at that position.
Before forming a union, RightTyper attempts to simplify the set of inferred types, counteracting the tendency of dynamic inference to produce large unions of concrete types. It does so by merging types subsumed by others (e.g., merging list[int] into list[int|str]) and by replacing groups of concrete types with an appropriate common supertype. To preserve semantic correctness, RightTyper verifies that the chosen supertype provides all methods and attributes shared by the replaced types. For numeric types, RightTyper follows Python’s “numeric tower”; for example, observing both int and float results in the simplified type float (pep484).
Simplifying for Human Use
Despite the simplifications above, annotations can still be complex. RightTyper offers options to trade precision for readability, such as collapsing repeated argument types (e.g., tuple[int, int, int] to tuple[int, ...]) and limiting nesting depth.
2.7. Handling Test-Borne Types
When a test suite drives execution, it often invokes production code with test-specific objects such as mocks. Such types should not appear in production annotations, as they do not reflect intended public APIs. RightTyper automatically detects test modules and excludes types defined in them from inferred annotations.
2.8. Annotating Source Code
The final step modifies source code to insert type annotations. By default, RightTyper only annotates unannotated program elements, preserving existing annotations; an option allows overwriting them when desired. When an inferred type requires a new import, RightTyper guards it using TYPE_CHECKING and emits string annotations (Figure 8). Static type checkers evaluate code under the equivalent of TYPE_CHECKING=True and resolve string annotations, so this approach avoids runtime import issues—particularly circular dependencies—while preserving full type checking.
Figure 8. RightTyper avoids runtime import errors: RightTyper guards added imports with TYPE_CHECKING and uses string annotations, avoiding runtime import issues while remaining compatible with static type checkers (excerpt from the pylint package).
3. Implementation
Our implementation of RightTyper comprises approximately 8,000 lines of Python code and is available as open-source software on GitHub (righttyper-github). This section describes key implementation details.
Instrumentation
RightTyper uses sys.monitoring events (PY_START, PY_RETURN, PY_YIELD, and PY_UNWIND) to capture function arguments, return values, yielded values, and exception-driven unwinding. Yielding temporarily suspends execution to produce a value before later resuming. Generator functions and coroutines (including async functions) both yield values, but coroutines may also receive values during execution; because sys.monitoring does not expose these received values, RightTyper supplements monitoring with AST-based instrumentation, intercepting code loading to inject the necessary hooks before execution.
Import Paths
Python’s runtime type information does not always yield usable import paths, particularly for built-in and C-implemented types. To address this, RightTyper constructs a mapping from type objects to import paths by scanning Python modules for type definitions. When multiple import paths exist, it applies heuristics to select a canonical one, prioritizing public exports (via __all__), avoiding internal names, and preferring shorter paths within the same package. This process produces more natural and consistent annotations across the codebase.
Sampling
Because most standard containers do not support efficient random sampling, RightTyper uses iterator-based sampling when necessary. To support statistically correct $\mathcal{O}(1)$ sampling for dictionaries, it optionally replaces dict with a custom implementation that maintains elements in an auxiliary list via AST-based instrumentation. This replacement is optional, as some programs rely on behavior specific to the standard dict.
Typing Iterators
Built-in iterators do not emit sys.monitoring events. To handle these cases, RightTyper uses Python’s garbage collection module to recover the underlying iterable (e.g., a container) and infer element types, falling back to Iterator[Any] when this is not possible. It identifies such iterators using a maintained list of expressions known to produce them.
Python Version Support
While it relies on sys.monitoring and therefore requires Python 3.12 or later to execute, it can emit annotations targeting Python 3.9 and later. Given the many extensions introduced since type annotations were first added to Python (typing-peps), RightTyper adapts its annotations to the target Python version, for example using int|str|None instead of Optional[Union[int, str]] for Python 3.10 and later.
Libraries
The implementation uses Python’s standard ast module for AST-based instrumentation and leverages the libcst concrete syntax tree library for static analysis and source-to-source transformation (libcst). It also uses typeshed_client to access the typeshed type stub repository (typeshed-client; typeshed), which is necessary because much of the standard library and many third-party packages lack inline type annotations.
4. Evaluation
Evaluating type inference is a challenging task. One of the difficulties lies in establishing a reliable ground truth. Some prior work derives ground truth from developer-written annotations, sometimes supplemented with types inferred by static analyzers such as pytype or Pyre (typilus; type4py). However, many projects contain type errors, as type checking is not yet a standard part of development workflows, and relying on developer-written annotations restricts evaluation to already-typed codebases (10.1145/3540250.3549114). Moreover, types inferred by static tools are only as reliable as the tools themselves: due to dynamic language features, these tools may conservatively over-approximate types or fail to infer them altogether (§5). Existing datasets also often omit support for code execution, as they target static and AI-based type inference methods (manytypes4py; typeevalpy; typybench).
To address these challenges, we evaluate RightTyper both on TypeEvalPy—a benchmark with an explicit ground truth, which we extend to support code execution—and on well-maintained real-world codebases, where inferred annotations can be compared against developer-written ones (typeevalpy).
Another difficulty lies in comparing inferred types. Prior work has often relied on string comparison of annotations (typewriter; manytypes4py; typilus; typeevalpy). But a program element containing a list[int], for example, can be validly annotated in multiple ways. In addition to trivial aliases like List[int] and less precise forms like list, it may also be appropriate to use protocol-based types such as Collection, Sequence, or Iterable. Because string comparison cannot recognize semantic relationships between types, as the nesting depth increases (e.g., dict[str, str|int|list[str|int]]), string match scores drop to 0 as it is unable to recognize partial matches (typybench). An effective comparison metric must be sufficiently nuanced to recognize when variants are appropriate and also indicate partial matches. Some prior work instead adopts a “correctness modulo type checker” approach to evaluation, measuring the quality of inferred types on whether they pass static type checking without errors (10.1145/3689783; typilus; 10.1145/3485488). However, doing so reinforces the often unwarranted assumption that the underlying code is correct. As Figure 9 illustrates, annotations can also effectively silence type checkers, inadvertently concealing bugs. For this evaluation, we use a modified version of the TypeSim type similarity metric proposed in prior work, which enables comparison with both established ground truth and developer-written annotations (§4.1). We do not use TypyBench’s TypeCheck metric, which counts mypy errors in annotated code. Dynamic approaches only annotate executed code, leaving unexecuted functions unannotated; when annotated code calls unannotated code (or vice versa), mypy reports type errors at these call sites even if the annotations themselves are correct. TypeCheck thus penalizes dynamic tools for incomplete coverage rather than annotation quality.
Figure 9. QuAC silences mypy: The code, previously shown in Figure 4, contains a function that, when invoked with two numbers, returns their sum, but when invoked with two strings, returns their concatenation. QuAC infers the SupportsAdd protocol, which prevents mypy from detecting the type mismatch error (middle). Executing the program yields a Python runtime error (bottom).
Our evaluation investigates the following questions:
RQ1::
Do type annotations generated by RightTyper improve upon those from prior approaches? (§4.2, §4.3)
RQ2::
How does RightTyper’s runtime overhead compare to that of other dynamic approaches? (§4.4)
Before addressing these questions, we describe the TypeSim metric used in our evaluation.
4.1. Metric
We compare annotations using a modified version of the TypeSim metric proposed by TypyBench (typybench). TypeSim provides a continuous measure of similarity between types based on the operations they support. It compares base types using the Jaccard index, defined as the ratio of shared attributes to the total number of attributes across both types:
$\operatorname{Jaccard}(t,t^{\prime})=\frac{|\operatorname{attrs}(t)\cap\operatorname{attrs}(t^{\prime})|}{|\operatorname{attrs}(t)\cup\operatorname{attrs}(t^{\prime})|}$
TypeSim operates recursively to support generic types. It handles unions by first computing an optimal matching between their elements and then comparing the matched elements.
The original implementation builds upon mypy and relies on the types it reports to compute similarity; this includes types inferred by mypy even in the absence of annotations, as well as types mapped to the Any wildcard when mypy cannot determine them (e.g., for dependencies lacking annotations) or abandons analysis due to earlier errors—which can either inflate or deflate TypeSim scores. We therefore reimplement TypeSim without relying on mypy. In addition, we correct an issue in the original implementation that compared only attributes declared directly on each type, yielding a TypeSim score of 0 between a type and its parent class.
We further extend TypeSim to support additional typing constructs—such as Callable, Literal, Type, and TypeAlias—which the original implementation ignored, and to better handle matching Any and object in generic parameters. Developers commonly use Any or object to avoid overly complex annotations (e.g., dict[str, Any] instead of dict[str, int|str|list[str]]).
4.2. [RQ1] Type Inference on TypeEvalPy
We address RQ1 by first evaluating RightTyper on TypeEvalPy, a benchmark suite of 153 short programs designed to evaluate type inference across a range of Python language features (typeevalpy). To make this possible, we add explicit calls to ensure code execution and adjust its assertions to target only program elements for which Python permits annotations (excluding, for example, lambda parameters), resulting in a total of 827 type assertions.
We then evaluate RightTyper on the modified suite. In doing so, we identify a number of discrepancies, all of which we determined to be errors in the benchmark’s ground-truth files. We corrected these errors and submitted the fixes to the benchmark authors, who accepted them. Finally, we add support for running QuAC, MonkeyType, pytype, and Type4Py as baselines. Given its apparent lack of maintenance, we do not include PyAnnotate in our evaluation. We also exclude Pyre, as its infer command is too limited to be useful: it cannot infer argument types from usage, return types from operations, or types for class methods. We execute MonkeyType without enabling sampling, allowing it to record every execution and thus achieve its best possible results. TypeEvalPy already supports prompting OpenAI models for type inference; we include results obtained by prompting the GPT-4o model using that implementation. Despite our best efforts to correct them, bugs in TypeT5 prevent us from including it in this comparison. RunTyper (10.1145/3771544) was also excluded as it requires a modified Python interpreter incompatible with standard installations.
Figure 10(a) presents the results in terms of TypeSim scores. With an overall 99.8% TypeSim, RightTyper achieves a near-perfect score, with only two cases failing to yield an exact match. In one case, the benchmark requests the type of an abstract method that does not execute. In the other case, RightTyper correctly types a zip iterator using the Iterator protocol—necessary to specify element types, since zip itself does not support them—but receives a lower score. This lower score occurs because TypeSim penalizes protocol-based annotations: protocols expose only the minimal set of required operations and therefore appear less similar to richer concrete types under the metric. GPT-4o, at 96.7%, also performs well, only incorrectly inferring “Unknown” or Any in a few cases. It is followed by pytype at 60.9%; it suffers from low coverage, consistent with prior work showing that its conservative static analysis leads to limited coverage (10.1145/3652153). MonkeyType and QuAC achieve 32.9% and 23.1%, respectively, in large part because they lack support for inferring variable types (Figure 11). QuAC operates by collecting sets of required attributes and using BM25—a term-frequency ranking function from information retrieval—to match these sets to candidate types (bm25). It is therefore plausible that the benchmark’s short programs—designed to facilitate manual ground-truth verification—provide limited context for this matching process, reducing QuAC’s effectiveness (10.1145/3689783). Similarly, Type4Py’s relatively low performance (24.7%) may be due to the benchmark programs likely representing out-of-distribution inputs for its model.
(a) TypeEvalPy benchmark (827 annotations)
(b) Real-world code (6,168 annotations)
Figure 10. [RQ1] RightTyper achieves the highest type inference accuracy: (a) On TypeEvalPy, RightTyper achieves 99.8% TypeSim, outperforming GPT-4o (96.7%). (b) On real-world code, RightTyper achieves 76.8% TypeSim, outperforming GPT-4o (60.6%) and MonkeyType (52.8%).
4.3. [RQ1] Type Inference on Real-World Code
Although the programs in TypeEvalPy provide broad coverage of Python language features, they are significantly simpler than, and differ markedly from, commonly deployed Python code. We therefore next evaluate RightTyper on the real-world Python packages pylint and black, which together contain over 6,000 annotations. We select these packages based on several criteria. First, both are among the most extensively annotated open-source Python projects, with annotations developed and refined by expert maintainers over many years of use with type checkers—making them ideal ground-truth sources. Second, they are complex, production-quality codebases that exercise diverse Python features including inheritance hierarchies, generic types, protocols, decorators, context managers, and AST manipulation. Third, they are widely used (pepy.tech reports billions of downloads) and well-maintained, reducing the risk of ground-truth errors. Finally, as command-line tools, we can drive them end to end with varied inputs, achieving broader coverage than test suites alone. As ground truth, we use the projects’ original type annotations; that is, we evaluate the semantic similarity of inferred annotations to developer-written ones using TypeSim.
Prior to evaluating each system, we remove all type annotations from the source code, except for special constructs such as dataclasses and NamedTuple where removing annotations would alter program semantics; in these cases, we replace the annotations with a type alias placeholder, allowing us to distinguish them from tool-inferred types during evaluation. We also remove any docstrings, as these can influence AI-based type inference; this ensures a level comparison with non-AI methods that cannot leverage natural language hints.
When evaluating RightTyper and MonkeyType, we execute each target package multiple times with varying options and inputs—keeping them identical across the two systems—and also run their test suites to improve coverage, thereby collecting typing information across a range of execution scenarios. As before (§4.2), we run MonkeyType without sampling, allowing it to record every execution and achieve its best possible results. Rather than having MonkeyType annotate the source code directly, we configure it to generate stub files, as in-source annotation causes MonkeyType to emit import statements that introduce circular dependencies. Applying MonkeyType and QuAC to these sources uncovered several bugs, including an assertion failure in QuAC and syntax errors in MonkeyType’s output; we corrected these issues. To include LLM-based results, we reimplement querying GPT-4o using TypyBench’s “single-file context” prompt, as its replication package does not provide an implementation.
Figure 10(b) presents the results. With an overall TypeSim score of 76.8%, RightTyper achieves the highest similarity to developer-annotated types. It is followed by GPT-4o and MonkeyType, which at 60.6% and 52.8% also show good performance, and at a considerable distance by Type4Py at 30.8%, QuAC at 26.7%, and pytype at 23.2%.
(a) Functions vs. variables breakdown
(b) All metrics on real-world code
Figure 11. [RQ1] Detailed results: (a) TypeSim scores by annotation category show RightTyper excels on both functions and variables, while QuAC and MonkeyType cannot infer variable types. (b) On real-world code, RightTyper achieves the highest scores across all metrics: TypeSim (76.8%), Exact Match (67.5%), and Coverage (89.8%).
Examining pytype’s annotations, we find that they often omit type arguments and exhibit low coverage, consistent with the results in §4.2. QuAC likewise suffers from low coverage, failing to infer types for nearly half of the function annotations—a limitation exacerbated by its lack of support for variable typing. Its annotations also often confuse nodes with similar attribute sets, a consequence of its BM25-based attribute-to-type mapping.
Type4Py’s annotations reveal frequent prediction errors, including a bias toward common types such as str, confusion between similarly named types from different libraries, omission of None for optional parameters, and difficulty handling project-specific types. These issues explain its low TypeSim score despite relatively high coverage.
Several factors affect MonkeyType’s score: it often infers overly concrete types where the original annotations use protocols (e.g., List[str] instead of Sequence[str]), misses some code paths, and lacks support for variable typing. GPT-4o performs relatively well on function annotations but poorly on variable annotations, which significantly lowers its overall score; it also frequently emits None for optional types or falls back to Any where more specific types were originally annotated.
RightTyper exhibits some similar limitations, including occasional overly concrete typing and missed coverage, the latter being expected given that it executes the same scenarios as MonkeyType. However, RightTyper achieves higher scores overall due to more complete language support—such as inferring variable types, handling keyword and variadic arguments, and inferring Callable argument types, all of which MonkeyType lacks—as well as its ability to better simplify inferred types, yielding cleaner annotations that more closely match the originals (Figure 11). The lower scores on variable annotations stem from variables captured only at function exit (missing intermediate types), unexecuted branches, and developers annotating with broader types than observed at runtime. Finally, in some cases RightTyper correctly emits annotations with type parameters (cf. §2.4), but TypeSim yields a zero score because it does not capture these forms of semantic equivalence.
Given RightTyper’s strong results on both TypeEvalPy and real-world code, we conclude that it substantially improves annotation quality relative to prior approaches.
[RQ1] Summary: With near-perfect inference on TypeEvalPy and the highest TypeSim on real-world code (77% vs. 61% for GPT-4o), RightTyper produces annotations of substantially higher quality than prior approaches—enabled by grounding types in actual runtime behavior and supporting the full range of Python language features.
4.4. [RQ2] Runtime Performance
[…]
To evaluate RightTyper’s runtime performance, we measure the time it requires to collect type information while running a suite of benchmark applications, normalized against each application’s baseline runtime. We perform the same measurements for two existing dynamic type inference tools, MonkeyType and PyAnnotate; given the limited number of available dynamic inference tools, we include PyAnnotate. We execute all tools with their default configurations.
We draw benchmarks from three sources. To represent real-world Python applications, we run the black code formatter on the large scikit-learn package and execute a Sudoku solver (sudoku-solver). We also run the test suites of two widely used Python packages, black and tornado. Finally, we include benchmarks drawn from Python’s benchmark suite (pyperformance).
We run all experiments using Python 3.12.8 on a 10-core 3.7GHz Core i9 system with 64GB of RAM and SSD storage, running Linux 6.5.6. With the system otherwise idle, we run each experiment five times and report the median execution time.
Table 1. [RQ2] RightTyper is fast: While Type4Py and QuAC are faster, RightTyper’s execution time on the real-world code from §4.3 remains within reasonable bounds, and faster than pytype and GPT-4o, in contrast to MonkeyType, which takes nearly four hours to complete.
Metric Type4Py QuAC MonkeyType pytype GPT-4o RightTyper
TypeSim 30.8% 26.7% 52.8% 23.2% 60.6% 76.8%
Time (h:mm) 0:03 0:04 3:58 0:11 0:12 0:07
Larger TypeSim are better; shorter times are better.
[…]
Table 1 shows the previously achieved TypeSim scores and running times. Type4Py and QuAC are the fastest, completing in 3 and 4 minutes, respectively, followed by RightTyper at 7 minutes and pytype at 11 minutes. GPT-4o completes in 12 minutes when issuing four prompts asynchronously; increased parallelism could likely reduce its execution time. MonkeyType is by far the slowest, requiring nearly four hours. We conclude that RightTyper is fast enough for practical use.
[…]
4.5. Threats to Validity
Internal Validity
Our evaluation depends on the correctness of our TypeSim reimplementation and the ground truth used. We mitigate this by releasing our implementation and TypeEvalPy corrections for community review. The choice of sampling parameters (k=5k{=}5, $\lambda{=}2\,\mathrm{Hz}$, $\tau{=}0.05$) may affect results; we use defaults that balance overhead and coverage based on preliminary experiments.
External Validity
Our real-world evaluation covers two Python packages; while both are development tools, they exercise diverse language features and contain annotations refined through years of production use. Results may differ for other domains (e.g., web frameworks or ML code), though RightTyper’s language-agnostic design imposes no domain-specific limitations. Additionally, when execution is driven by test suites, RightTyper’s coverage depends on test quality: codebases with limited test coverage may yield incomplete annotations, while tests that exercise unintended behavior may result in annotations that reflect that behavior.
Construct Validity
TypeSim measures structural similarity via shared attributes, which may not fully capture semantic equivalence—it penalizes protocol-based annotations and does not recognize type parameter equivalence. Developer-written annotations serve as ground truth but may contain errors or stylistic choices rather than the most precise types.
5. Related Work
There has been substantial research on type inference for dynamically typed languages, particularly JavaScript, Ruby, and Python. This section concentrates primarily on prior work targeting Python.
Static type inference
Static approaches analyze program structure without executing it, relying on techniques such as abstract interpretation, data-flow analysis, and type constraint resolution to infer types (maia2012static; pysonar; 10.1145/2661088.2661101; 10.1145/3551349.3561150; DBLP:conf/cav/HassanUE018; 10.1145/3689783; monat_et_al:LIPIcs.ECOOP.2020.17).
Static type checkers such as mypy, pyright, pyre, and pytype must also perform type inference to enable type checking, though their inference may be limited, primarily focused on how types from existing annotations propagate through the code or are affected by operations (mypy; pyright; pyre; pytype). Among these tools, only pytype and pyre support adding type annotations.
While typically sound by design, static approaches are constrained by dynamic language features such as runtime type changes, reflection, monkey patching, and dynamic code generation, and often support only a subset of their target languages (10.1145/1806596.1806598; hityper; 10.1007/978-3-642-03237-0_17; DBLP:conf/cav/HassanUE018). Dynamic features can also lead static systems to conservatively over-approximate, inferring overly permissive types (ruby-dynamic; typewriter). In contrast to static approaches, RightTyper produces precise annotations grounded in actual runtime behavior for all executed code paths.
AI-based type inference
These approaches estimate the likely types of program elements from features extracted from the source code, often including natural language elements. Xu et al. apply a probabilistic graphical model to predict variable types based on hints such as attribute accesses and variable names (10.1145/2950290.2950343). TypeWriter applies a neural network to both code and natural language information and verifies the predictions using a type checker (typewriter). Typilus and Type4Py leverage deep similarity learning to avoid using a fixed type vocabulary (typilus; type4py). PYInfer uses a neural network to infer variable types from the surrounding source code context (pyinfer). HiTyper combines static and machine learning inference by only applying the machine learning model where its type constraints resolution gets stuck (hityper). Abdelaziz et al. combine mining docstrings and approximate duck subtyping to infer function return types (abdelaziz-et-al-2022; duck-typing). TypeT5 fine-tunes CodeT5 for type inference (typet5; codet5). DLInfer extracts program slices for variables and applies a sequence model to predict function argument types (10.1109/ICSE48619.2023.00170). TypeGen prompts ChatGPT with chain-of-thought instructions to infer types (typegen; chain-of-thought). TIGER ranks generative model typing candidates with a similarity model (tiger). DeMinify, primarily targeting minified code recovery, uses dual-task learning between variable name and type prediction (deminify). RunTyper combines dynamic analysis with deep neural networks to validate and refine type predictions; it employs a modified Python interpreter to gather runtime type information (10.1145/3771544).
While better able to cope with Python’s dynamic nature than purely static approaches, AI-based methods are inherently unsound: their predicted types may not reflect actual program behavior. By combining AI-based and static or dynamic inference, approaches such as TypeWriter, HiTyper, Typilus, and RunTyper can reject incorrect types—but risk leaving more of the code unannotated (typewriter; hityper). Furthermore, these approaches often only support a limited, small type vocabulary and perform poorly on rare types, making them less effective when encountering custom or evolving types (type4py). Unlike AI-based approaches, RightTyper produces annotations grounded in actual runtime behavior.
Dynamic type inference
Dynamic approaches track types encountered during program execution. Tools such as MonkeyType (Instagram), PyAnnotate (Dropbox, developed by Python author Guido van Rossum), and RightTyper employ this strategy (monkeytype; pyannotate). A major limitation of prior dynamic approaches is the runtime overhead they impose. Although both MonkeyType and PyAnnotate support sampling of function calls to mitigate overhead, their designs require some instrumentation to remain continuously active, incurring nontrivial cost. MonkeyType implements Bernoulli random sampling (disabled by default), but even when enabled, this requires instrumentation on every call to decide whether to sample, and MonkeyType logs all collected data to a SQLite database. Moreover, for each sampled call, MonkeyType recursively and exhaustively scans every element of every container argument—resulting in severe performance degradation for large containers and, in some cases, causing execution to abort due to timeouts or resource exhaustion. In contrast, RightTyper uses a Poisson process: instrumentation is disabled except during brief capture windows, so most calls incur zero overhead. Because capture windows occur at random times, every call has an equal probability of being observed, ensuring representative sampling without systematic bias. PyAnnotate’s deterministic strategy also avoids per-call decisions but can introduce substantial bias; RightTyper uses a Good–Turing estimator for containers, which is both efficient and statistically principled.
Both MonkeyType and PyAnnotate rely on Python’s runtime type attributes, which are not always reliable. MonkeyType prunes invalid type names, but this can leave code unannotated; PyAnnotate performs no validation and may emit annotations causing runtime errors. RightTyper instead maps type objects to valid, public names.
Additionally, MonkeyType and PyAnnotate lack support for important language features: inferring variable types, handling keyword and variadic arguments, typing most non-container generics, and correctly typing inherited methods. When types vary across executions, they construct simple unions that can be incorrect or overly permissive. In contrast, RightTyper supports these features, recognizes recurring type patterns, and applies a suite of context-aware simplification strategies where appropriate (§2.22.6). As far as we are aware, no typing tool other than RightTyper supports annotating array dimensions (§2.3).
6. Conclusion
This paper introduces RightTyper, a novel hybrid type inference approach for Python that produces accurate and precise type annotations. By combining dynamic analysis with static analysis, type aggregation, and name resolution, RightTyper generates annotations of substantially higher quality than those produced by prior approaches. Its principled, statistically guided adaptive sampling enables it to observe representative program behavior while maintaining low runtime overhead. Together, these characteristics make RightTyper a practical and effective tool for annotating real-world Python programs, enabling developers to benefit from static type checking with minimal manual effort.
Future work includes extending RightTyper to additional language features and improving support for already annotated codebases.
Data Availability
RightTyper is available as source code at https://github.com/RightTyper/RightTyper.